# Mixture of Experts based Multi-task Supervise Learning from Crowds

### v1

@@ 5 Experiments
  • Text Dumitrache et al., (2018): This dataset comprises 1594 sentences extracted from the CrowdTruth corpus, categorized into 13 groups. The dataset includes 14,228 labels provided by 154 workers. Each sentence is represented by 768-dimensional features extracted using a pre-trained BERT model.
  • Music Rodrigues et al., (2014): This dataset consists of 700 music compositions, each with a duration of 30 seconds, and categorized into 10 groups. It includes 2,945 labels provided by 44 workers. Each music composition is represented by 124-dimensional features extracted using the Marsyas Rodrigues et al., (2013) music retrieval tool.
@@
To accommodate the feature scales of the three experimental datasets, our model’s architecture varies accordingly. For the LableMe dataset, our model employs 16 expert modules, each comprising 3 fully connected ReLU layers, with a final layer output dimension of 32. For the Text and Music datasets, we utilize 10 expert modules. Each module consists of 3 and 2 fully connected ReLU layers, with output dimensions of 32 and 16, respectively. We adopt the settings from the Max-MIG, CoNAL, and CrowdAR truth inference methods, we adopt the settings from their source code for the LableMe and Music datasets. Since there is no source code available for the Text dataset, we adopt the settings used in the LableMe dataset. Regarding the TDG4Crowd data filling algorithm, we utilize the settings from its source code. The remaining methods do not use deep network structures and rely on default settings.
5.1 Evaluation of Oracle Worker Finding (MMLC-owf)
@@
Table 1: Accuracy of Truth Inference Methods on Three Crowdsourced Datasets.
@@
Figure 2: Accuracy of MMLC-owf with Various Clustering Methods.
@@
Main Result: Our method, MMLC-owf, was evaluated alongside seven other methods through five rounds experiments. The average accuracy results are shown in Tab. 1. In our method, we utilized kernel density estimation (KDE) to compute the projection vector of the oracle worker in the worker spectral space. Our method, MMLC-owf, achieved the highest accuracy in the Text and Music datasets. In the LableMe dataset, it ranked second, with only a 0.4% difference from the top-performing CrowdAR method. Deep learning-based methods typically produce better results when analyzing datasets with high-dimensional item features like LableMe. In datasets with fewer features, the advantage of deep learning methods was not significant.
@@
Impact of Redundancy: We examine how varying levels of redundancy affect the accuracy of our method. Due to the varying redundancy of data items, a maximum redundancy parameter $R$ is set. We randomly keep $R$ labels for items with more than $R$ labels and discard the rest. This process generates a dataset with a maximum redundancy $R$. By conducting five repeated experiments and averaging the accuracy and standard deviation, the results are shown in Fig. 3. As the average number of worker responses increases, all methods show an upward trend in their results. The analysis of various redundancy levels across the datasets indicates that higher redundancy levels are more advantageous for our method. Our method can effectively utilize worker behavior descriptions on datasets with higher redundancy but may face underfitting on datasets with lower redundancy.
@@
Clustering Methods in Oracle Worker Finding: Our method, MMLC-owf, utilizes a clustering method to determine the center of the distribution of the projection vector of workers in the worker spectral space as the projection vector of the oracle worker. Here, we examine how different clustering methods affect truth inference outcomes. We compare three clustering methods: kernel density estimation (KDE), Mean, and Median, as well as their worker-weighted variants: KDE-W, Mean-W, and Median-W. Worker weights are calculated based on the proportion of items answered by each worker relative to the total number of items, considering data imbalance. In addition, we optimize the projection vector in the worker spectral space using the ground truth of the items as an upper-bound method for clustering. The parameters of the expert modules and output modules are fixed in the pre-trained MMLC model, referred to as “Truth.” By conducting five repeated experiments and averaging the results, as shown in Fig. 2. For example, in the LableMe dataset, the oracle worker’s projection vector is derived using KDE, KDE-W, Median, Median-W, Mean, and Mean-W. The MMLC-owf uses the oracle worker to generate ground truth. The quality of the ground truth obtained by the following five methods in each dataset is very similar, but the oracle worker generated using KDE produces the highest quality ground truth in each dataset. The “Truth" method, which represents the theoretical upper limit with clustering methods, achieved accuracy rates of 96.32%, 91.25%, and 92.97% on the three datasets respectively. The quality of the ground truth generated by oracle workers using six clustering methods still slightly deviates from theoretical upper limits. This implies that the MMLC-owf method can provide high-quality ground truth by optimally projecting the worker spectral space, approaching the theoretical upper limit. The model has strong expressive ability, with a small gap between the projected spectral space and the ground truth. There is potential to enhance MMLC-owf by choosing a more effective clustering method.
(a) LabelMe
(b) Text
(c) Music
Figure 4: Worker Scatter Plot in Worker Spectral Space.
@@
Worker Distribution in Worker Spectral Space: We assume that each worker is an oracle worker with random errors in their expression. The center of the distribution of workers projected onto the spectral space corresponds to the projection vector of the oracle worker. To validate this assumption, we employed the IOSMAP dimensionality reduction method to reduce the worker projection vectors obtained from the MMLC model to 2D, resulting in the scatter plot shown in Fig. 4. We calculated the accuracy of each worker on the dataset, where the closer the point’s color on the graph is to green, the worker’s accuracy is higher. The closer the point’s color is to red, the lower the worker’s accuracy. The plot also shows the projection obtained by the KDE method for the oracle worker, represented by blue asterisks. From the distribution of worker projections, although the shapes of the distributions differ across datasets, there is a noticeable trend where workers with higher accuracy tend to cluster closer to the projection of the oracle worker. This observation demonstrates a clear tendency towards aggregation and provides some degree of confirmation for the validity of our assumption.
5.2 Evaluation of Data Filling (MMLC-df)
@@
Main Result: We compared MMLC-df with three filling methods: G_MV, G_IRT, and TDG4Crowd. We used the filled data with eight truth inference methods from the previous section to infer the ground truth. We used three real datasets and applied eight truth value inference methods to infer the ground truth, resulting in a total of 24 scenarios. We conducted five rounds of experiments, and the mean and variance of all ground truth accuracy are presented in Tab. 2. It can be observed that our MMLC-df framework achieves enhanced performance compared to the original data in 100% of the scenarios, with 79.2% of the scenarios achieving optimal enhancement. On the other hand, G_MV, G_IRT, and TDG4Crowd achieve enhanced results in 50%, 62.5%, and 75% of the scenarios, respectively. In terms of the enhancement magnitude, our method performs the best on the Text dataset. While other methods may achieve better results in certain scenarios for other datasets, their performance is unstable, and there are cases where the results deteriorate. This indicates that our MMLC-df framework demonstrates good stability and consistent performance.
Impact of Data Filling’s Density: Our MMLC-df framework leverages the sparsity of crowdsourced data for data filling. To clarify, we define the data density of non-empty crowdsourced data as $d_{\mathcal{D}}\in(0,1]$ and $d_{\mathcal{D}}=\frac{|\mathcal{D}|}{|\mathcal{W}|\times|\mathcal{X}|}$. The data densities of the three original datasets LabelMe, Text, and Music are 0.0431, 0.0579, and 0.0956, respectively. The original data seems sparse. We gradually fill the data until reaching a data density to 1 for the analysis of its impact of data density on the results. We set a threshold for the number of items to be filled, denoted as $n_{interval}$. For workers with items exceeding this threshold, we replace their labeled items with predicted values. By adjusting the threshold from large to small, we gradually fill the data until all workers have completed their items, achieving a data density of 1. Due to the large amount of filled data, deep learning methods can be time-consuming. The accuracy obtained by various methods shows a similar trend to the density transformation. Therefore, we conducted this experiment using only statistical machine learning methods. The experiment is repeated five rounds, and the average accuracy and standard deviation are shown in Fig. 5. The trends are generally consistent across all methods, but the variations differ significantly among different datasets. In Text dataset, as density increases, the algorithm’s accuracy stabilizes rapidly and then reaches a plateau. In the LableMe dataset, accuracy fluctuates significantly as density increases. Higher density often improves accuracy. In Music dataset, as density increases, accuracy initially fluctuates rapidly before stabilizing. The Text data filling performs the best, likely due to the larger scale of this dataset compared to the other two, resulting in a more significant impact.
@@
6 Conclusion
This paper introduces a novel crowdsourced learning paradigm called MLC.Within this paradigm, we propose a feature-level worker behavior model called MMLC. Based on this model, we develop two truth inference methods: MMLC-owf, which uses oracle worker finding, and MMLC-df, a truth inference framework based on crowdsourced data filling. Experimental results demonstrate the superior performance of MMLC-owf compared to other methods. Furthermore, we assess the theoretical upper performance limit of the MMLC-owf method, demonstrating its potential to enhance clustering method selection and validate its strong performance. The experiments also validate the effectiveness and stability of the MMLC-df framework in enhancing truth inference methods through crowdsourced data filling. Furthermore, we observed that our model exhibited better performance on datasets with a higher number of annotations per worker. On real crowdsourcing platforms, workers continuously engage in annotation tasks, resulting in an increasing average number of annotations per worker. Consequently, our model holds significant practical value for real-world applications.

### v2

@@ Introduction
We identified three representative datasets that exemplify different types of crowdsourcing scenarios and data characteristics. LableMe (Rodrigues and Pereira 2018; Russell et al. 2008): This dataset consists of 1000 images categorized into 8 classes, with a total of 2547 labels provided by 59 workers. Each image is represented by 8192-dimensional features extracted using a pre-trained VGG-16 model. Text (Dumitrache, Aroyo, and Welty 2018): This dataset comprises 1594 sentences extracted from the CrowdTruth corpus, categorized into 13 groups. The dataset includes 14,228 labels provided by 154 workers. Each sentence is represented by 768-dimensional features extracted using a pre-trained BERT model. Music (Rodrigues, Pereira, and Ribeiro 2014): This dataset consists of 700 music compositions, each with a duration of 30 seconds, and categorized into 10 groups. It includes 2,945 labels provided by 44 workers. Each music composition is represented by 124-dimensional features extracted using the Marsyas (Rodrigues, Pereira, and Ribeiro 2013) music retrieval tool.
@@
To accommodate the feature scales of the three experimental datasets, our model’s architecture varies accordingly. For the LableMe dataset, our model employs 16 expert modules, each comprising 3 fully connected ReLU layers, with a final layer output dimension of 32. For the Text and Music datasets, we utilize 10 expert modules. Each module consists of 3 and 2 fully connected ReLU layers, with output dimensions of 32 and 16, respectively. We adopt the settings from the learning-from-crowds methods Max-MIG, CoNAL, and CrowdAR, we adopt the settings from their source code for the LableMe and Music datasets. Since there is no source code available for the Text dataset, we adopt the settings used in the LableMe dataset. The hyperparameters mainly follow the expert configuration from Google’s MMoE model(Ma et al. 2018) and the classifier setup of CrowdAR. Regarding the TDG4Crowd data filling algorithm, we utilize the settings from its source code. The remaining methods do not use deep network structures and rely on default settings.
@@
Table 1: Accuracy of Learning-from-Crowds Methods on Three Crowdsourced Datasets.
@@
Figure 3: Accuracy of MMLC-owf with Various Clustering Methods on Three Crowdsourced Datasets.
Evaluation of Oracle Worker Finding(MMLC-owf)
Main Result:
@@
Our method, MMLC-owf, was evaluated alongside seven other methods through five rounds experiments. The average accuracy results are shown in Tab. 1. In our method, we utilized kernel density estimation (KDE) to compute the projection vector of the oracle worker in the worker spectral space. Our method, MMLC-owf, achieved the highest accuracy in the Text and Music datasets. In the LableMe dataset, it ranked second, with only a 0.4% difference from the top-performing CrowdAR method. Deep learning-based methods typically produce better results when analyzing datasets with high-dimensional item features like LableMe. In datasets with fewer features, the advantage of deep learning methods was not significant.
Impact of Redundancy:
@@
We examine how varying levels of redundancy affect the accuracy of our method. Due to the varying redundancy of data items, a maximum redundancy parameter $R$ is set. We randomly keep $R$ labels for items with more than $R$ labels and discard the rest. This process generates a dataset with a maximum redundancy $R$. By conducting five repeated experiments and averaging the accuracy and standard deviation, the results are shown in Fig. 2. As the average number of worker responses increases, all methods show an upward trend in their results. The analysis of various redundancy levels across the datasets indicates that higher redundancy levels are more advantageous for our method. Our method can effectively utilize worker behavior descriptions on datasets with higher redundancy but may face underfitting on datasets with lower redundancy.
Clustering Methods in Oracle Worker Finding:
@@
Our method, MMLC-owf, utilizes a clustering method to determine the center of the distribution of the projection vector of workers in the worker spectral space as the projection vector of the oracle worker. Here, we examine how different clustering methods affect learning-from-crowds outcomes. We compare three clustering methods: kernel density estimation (KDE), Mean, and Median, as well as their worker-weighted variants: KDE-W, Mean-W, and Median-W. Worker weights are calculated based on the proportion of items answered by each worker relative to the total number of items, considering data imbalance. In addition, The parameters of the expert modules and output modules are fixed and we optimize the projection vector in the worker spectral space using the ground truth of the items as the best performance of our model for clustering. This result is referred to as “Truth.” By conducting five repeated experiments and averaging the results, as shown in Fig. 3. For example, in the LableMe dataset, the oracle worker’s projection vector is derived using KDE, KDE-W, Median, Median-W, Mean, and Mean-W. The MMLC-owf uses the oracle worker to generate ground truth. The quality of the ground truth obtained by the following five methods in each dataset is very similar, but the oracle worker generated using KDE produces the highest quality ground truth in each dataset. The “Truth” method, which represents the theoretical upper limit with clustering methods, achieved accuracy rates of 96.32%, 91.25%, and 92.97% on the three datasets respectively.The quality of the ground truth generated by oracle workers using six clustering methods still slightly deviates from theoretical upper limits. This implies that the MMLC-owf method can provide high-quality ground truth by optimally projecting the worker spectral space, approaching the theoretical upper limit. The model has strong expressive ability, with a small gap between the projected spectral space and the ground truth. There is potential to enhance MMLC-owf by choosing a more effective clustering method.
Worker Distribution in Worker Spectral Space:
@@
We assume that each worker is an oracle worker with random errors in their expression. The center of the distribution of workers projected onto the spectral space corresponds to the projection vector of the oracle worker. To validate this assumption, we employed the IOSMAP dimensionality reduction method to reduce the worker projection vectors obtained from the MMLC model to 2D, resulting in the scatter plot shown in Fig. LABEL:fig:scatter. We calculated the accuracy of each worker on the dataset, where the closer the point’s color on the graph is to green, the worker’s accuracy is higher. The closer the point’s color is to red, the lower the worker’s accuracy. The plot also shows the projection obtained by the KDE method for the oracle worker, represented by blue asterisks. From the distribution of worker projections, although the shapes of the distributions differ across datasets, there is a noticeable trend where workers with higher accuracy tend to cluster closer to the projection of the oracle worker. This observation demonstrates a clear tendency towards aggregation and provides some degree of confirmation for the validity of our assumption.
@@
Evaluation of Data Filling (MMLC-df)
Main Results:
We compared MMLC-df with three filling methods: G_MV, G_IRT, and TDG4Crowd. We used the filled data with eight learning-from-crowds methods from the previous section to infer the ground truth. We used three real datasets and applied eight truth value inference methods to infer the ground truth, resulting in a total of 24 scenarios. We conducted five rounds of experiments, and the mean and variance of all ground truth accuracy are presented in Tab. 2. It can be observed that our MMLC-df framework achieves enhanced performance compared to the original data in 100% of the scenarios, with 79.2% of the scenarios achieving best enhancement. On the other hand, G_MV, G_IRT, and TDG4Crowd achieve enhanced results in 50%, 62.5%, and 75% of the scenarios, respectively. In terms of the enhancement magnitude, our method performs the best on the Text dataset. While other methods may achieve better results in certain scenarios for other datasets, their performance is unstable, and there are cases where the results deteriorate. This indicates that our MMLC-df framework demonstrates good stability and consistent performance.
Impact of Data Filling’s Density:
Our MMLC-df framework leverages the sparsity of crowdsourced data for data filling. To clarify, we define the data density of non-empty crowdsourced data as $d_{\mathcal{D}}\in(0,1]$ and $d_{\mathcal{D}}=\frac{|\mathcal{D}|}{|\mathcal{W}|\times|\mathcal{X}|}$. The data densities of the three original datasets LabelMe, Text, and Music are 0.0431, 0.0579, and 0.0956, respectively. The original data seems sparse. We gradually fill the data until reaching a data density to 1 for the analysis of its impact of data density on the results. We set a threshold for the number of items to be filled, denoted as $n_{interval}$. For workers with items exceeding this threshold, we replace their labeled items with predicted values. By adjusting the threshold from large to small, we gradually fill the data until all workers have completed their items, achieving a data density of 1. Due to the large amount of filled data, deep learning methods can be time consuming. The accuracies obtained by various methods show a similar trend to the density transformation. Therefore, we conducted this experiment using only statistical machine learning methods. The experiment is repeated in five rounds, and the average accuracy and standard deviation are shown in Fig. 6. The trends are generally consistent across all methods, but the variations differ significantly among different datasets. In Text dataset, as density increases, the algorithm’s accuracy stabilizes rapidly and then reaches a plateau. In the LableMe dataset, accuracy fluctuates significantly as density increases. Higher density often improves accuracy. In Music dataset, as density increases, accuracy initially fluctuates rapidly before stabilizing. The Text data filling performs the best, likely due to the larger scale of this dataset compared to the other two, resulting in a more significant impact.
Conclusion
This paper introduces a novel crowd-learning paradigm called MLC. Within this paradigm, we propose a feature-level worker behavior model called MMLC. Based on this model, we develop two learning-from-crowds methods: MMLC-owf, which uses oracle worker finding, and a framework MMLC-df based on data filling. Experimental results demonstrate the superior performance of MMLC-owf compared to other methods. Furthermore, we assess the theoretical upper performance limit of the MMLC-owf method, demonstrating its potential to enhance clustering method selection and validate its strong performance. The experiments also validate the effectiveness and stability of the MMLC-df framework in enhancing learning-from-crowds methods through data filling. Furthermore, we observed that our model exhibited better performance on datasets with a higher number of annotations per worker.
Acknowledgments
This research has been supported by the Natural Science Foundation of Zhejiang Province, China (Grant No. LQ22F020002, LZ22F020008, and LQ24F020003), the National Natural Science Foundation of China under grant 61976187. Besides, the authors want to thank the anonymous reviewers for the helpful comments and suggestions to improve this paper.